Calculate the number of moles of CO2 formed when 4.3 moles of C3H8a reactive oxygen in the combustion reaction
1
Expert's answer
2013-02-15T05:36:36-0500
Calculate the number of moles of CO2 formed when 4.3 moles of C3H8a reactive
oxygen in the combustion reaction

C3H8 + 5O2 = 3CO2 + 4H2O

if 4.3 moles of propane react, then:
n(CO2) = 3n(C3H8)
n(CO2) = 3*4.3mol = 12.9 mol